# Quench Spectroscopy for Dissipative and Non-Hermitian Quantum Lattice Models

### v1

@@ 8 Conclusion
In this work we have discussed the quench spectroscopy of the open Bose-Hubbard chain confined in the superfluid phase where the dissipation is characterised by on-site two-body losses and of the non-Hermitian transverse-field Ising chain confined in the paramagnetic phase; both from a theoretical point of view.
@@
Regarding the dissipative Bose-Hubbard chain, we have first introduced the theory based on an EoM approach. Then, we benchmarked the latter with a quasiparticle theory for two distinct connected equal-time correlation functions while considering the isolated Bose-Hubbard chain. The latter functions are the $G_{1}$ one-body correlation function as well as the $G_{2}$ density-density correlation function giving insight regarding the phase and density fluctuations of the quantum model respectively. Then, we investigated the double quench dynamics of the Bose-Hubbard chain where both the repulsive interactions and the dissipations are quenched. Finally, we studied its quench spectroscopy by calculating the quench spectral function associated to $G_{1}$ and $G_{2}$. Concerning the non-Hermitian transverse-field Ising chain, we considered the $z$ polarized phase and its associated quench dynamics by considering a equal-time spin-spin correlation function in the aim to apply the quench spectroscopy method to the latter.
This paper unraveled new theoretical results regarding the double quench dynamics of the Bose-Hubbard chain as well as the quench spectroscopy of the dissipative Bose-Hubbard chain which can be confirmed experimentally using quantum simulators based on ultracold atoms. The applicability of the quench spectroscopy method is also extended to non-Hermitian quantum lattice models. To summarize, in the framework of weak sudden global quantum quenches, this paper paves the way to the possibility of applying the quench spectroscopy method to dissipative and non-Hermitian quantum lattice models.
@@
As an extension of the presented research work, the generalizability of the quench spectroscopy approach is an interesting research topic. The latter is likely to be generalized to dissipative fermionic lattice models, dissipative spin lattice models as well as to their reformulation in the continuum and for a higher dimensionality of the lattice or the space respectively. The quench spectroscopy method is also expected to be reliable for dissipative quantum systems involving long-range interactions and for other Lindblad jump operators, i.e. other kind of loss processes.
Acknowledgements
@@
The Author wishes to acknowledge M. Cheneau for fruitful discussions concerning the experimental feasibility of the quench spectroscopy method using cold-atom-based quantum simulators. The Author is also grateful to L. Sanchez-Palencia for useful comments regarding the manuscript.
Funding information
This research was supported by Region Ile-de-France in the framework of DIM QuanTip and by the ANR project LOQUST ANR-23-CE47-0006-02.

### v2

@@ VI Conclusion
In this work we have discussed the quench spectroscopy method for dissipative, isolated non-Hermitian and Hermitian quantum lattice models theoretically. For dissipative and isolated non-Hermitian quantum systems, we considered the case study of the open Bose-Hubbard chain confined in the superfluid phase where the dissipation is characterized by on-site two-body losses and the non-Hermitian transverse-field Ising chain confined in the paramagnetic phase respectively. For isolated Hermitian quantum lattice models, we used the Hermitian version of the latter quantum system.
@@
Regarding the dissipative Bose-Hubbard chain, we have first introduced the theoretical method based on the equation-of-motion approach. We then benchmarked the latter with the quasiparticle theory for two distinct equal-time connected correlation functions while considering the isolated Bose-Hubbard chain. The latter functions are $G_{1}$ the one-body correlation function as well as $G_{2}$ the density-density correlation function giving insight on the phase and density fluctuations of the quantum system respectively. We then investigated the double quench dynamics of the Bose-Hubbard chain where both the repulsive interactions and the dissipations are quenched. We finally studied the quench spectroscopy of this quantum model by calculating the quench spectral function associated to the one-body and density-density correlation functions.
Concerning the non-Hermitian transverse-field Ising chain confined in the paramagnetic phase, its quench dynamics has been investigated using the equation-of-motion approach and by considering an equal-time spin-spin correlation function in order to apply the quench spectroscopy method.
For the Hermitian case, we considered the Hermitian version of the previous model while considering the same gapped quantum phase. The advantage of this quantum model is twofold: the possibility to express the Hamiltonian in a quadratic fermionic and bosonic form and the possibility to rely on the equation-of-motion and quasiparticle theoretical approaches. This has permitted to unveil that for generic isolated and Hermitian quantum lattice models displaying a quadratic form, the quench spectroscopy method is suitable and reliable. To summarize, in the framework of weak sudden global quantum quenches, this paper paves the way to the possibility of applying the quench spectroscopy method to dissipative, isolated non-Hermitian and isolated Hermitian quadratic quantum lattice models.
@@
As an extension of the presented research work, the generalizability of the quench spectroscopy method is an interesting research topic. The latter is likely to be extended to dissipative fermionic lattice models, dissipative spin lattice models and to their reformulation in the continuum and for a higher dimensionality of the lattice or the space respectively. This new spectroscopy method is also expected to be reliable for dissipative quantum systems involving long-range interactions and for other Lindblad jump operators, i.e. other kind of loss processes.
While considering large timescales, another possible extension would be to investigate the applicability of the quench spectroscopy method while using quantum computing platforms. In a first time, this can be done at the architecture level using a digital twin of the quantum computer. In a second time, provided that the latter investigation is conclusive, one may consider to study this topic at the experimental level. Since the quench spectroscopy method requires a relatively large quantum system, i.e. many logical qubits, suitable experimental platforms are neutral-atom-based quantum computers and trapped ion quantum computers; the latter being derived from well-known quantum simulators.
Acknowledgments
@@
The Author acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101002955 – CONQUER), from the Région Île-de-France in the framework of DIM QuanTiP and from the ANR project LOQUST ANR-23-CE47-0006-02. The Author wishes to acknowledge M. Cheneau for fruitful discussions concerning the experimental feasibility of the quench spectroscopy method using cold-atom-based quantum simulators. The Author is also grateful to L. Sanchez-Palencia for useful comments regarding the manuscript. Finally, the Author would like to thank M. Schirò and L. Mazza for useful discussions regarding the dissipative quench dynamics of the Bose-Hubbard model.
Appendix A The non-connected equal-time density-density correlation function in the SF-mean-field regime and the condensate density in the non-interacting case
In what follows, we discuss the analytical expression of $g_{2}(R,t)$ and $g_{1}(R,t)$ the non-connected density-density and non-connected one-body correlation functions. As previously mentioned, the latter can be found using the set of EoMs defined at Eq. (8). As a reminder, $n(t)=N(t)/L=\langle\hat{N}\rangle_{t}/L$ refers to the time-dependent density (or equivalently the time-dependent filling) with $\hat{N}=\sum_{R}\hat{n}_{R}$ denoting the total occupation number operator; $g_{2}(R,t)=\langle\hat{n}_{R}\hat{n}_{0}\rangle_{t}$ and $g_{1}(R,t)=\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{t}$. We remind the reader that the latter correlation functions respectively permit to characterize the density and phase fluctuations and can be measured in experimental setups based on ultra-cold atoms using fluorescence microscopy imaging and time-of-flight techniques cheneau2012 ; trotzky2012 ; langen2013 ; geiger2014 . To derive their analytical expression, a same theoretical approach is performed namely: (i) we first express them in the reciprocal space (ii) we then consider a decoupling of the condensate mode $k=0$ from the finite modes $k\neq 0$ (iii) we simplify the expression using the standard mean-field approximation, i.e. $N(t)\gg\sum_{k\neq 0}G_{k}(t)$, as well as the product state approximation in momentum space $\langle\hat{n}_{0}^{2}\rangle_{t}=\langle\hat{n}_{0}\rangle_{t}^{2}$. The two previous approximations are valid provided that the SF-mean-field regime is considered as well as small observation times. This mathematical procedure leads to:
$g_{2}(R,t)=$ $\leavevmode\nobreak\ \frac{1}{L^{2}}\Big{\{}G_{0}(t)^{2}+\sum_{k\neq 0}\left(e^{ikR}F_{0}(t)F_{k}(t)^{*}+\mathrm{h.c}\right)+e^{-ikR}G_{0}(t)+e^{ikR}G_{k}(t)+2[1+\cos(kR)]G_{0}(t)G_{k}(t)\Big{\}},$ (78a)
$g_{1}(R,t)=$ $\leavevmode\nobreak\ \frac{1}{L}\Big{[}G_{0}(t)+\sum_{k\neq 0}\cos(kR)G_{k}(t)\Big{]}.$ (78b)
Note that this theoretical expression for $g_{1}$ is consistent with our previous findings, see Eq. (16) and the discussion beforehand. In what follows, we explain how to simplify the previous analytical form of $g_{2}$. To do so, we rely on the following approximations: $\operatorname{Re}(F_{0}(t))\gg|\operatorname{Im}(F_{0}(t))|$, $|\sum_{k\neq 0}\operatorname{Re}(F_{k}(t))|\gg|\sum_{k\neq 0}\operatorname{Im}(F_{k}(t))|$ and $N(t)\gg\sum_{k\neq 0}G_{k}(t)$ despresnew . The latter are deduced from the EoMs at Eq. (8) and valid in the SF-mean-field regime while considering the thermodynamic limit, i.e. $L\rightarrow+\infty$, and small observations times. This finally leads to Eq. (21). Note also that the condensate density for the non-interacting case, i.e. $U=0$, given at Ref. despresnew can be recovered using the previous expression of $g_{1}$ and $g_{2}$. Indeed, for $U=0$, we have $N(t)=G_{0}(t)$ and using the standard mean-field approximation, i.e. $N(t)\gg\sum_{k\neq 0}G_{k}(t)$, we can significantly simplify the expression of $g_{1}$ and $g_{2}$ given at Eq. (78). It yields:
$g_{2}(R,t)=n(t)\Big{[}n(t)+\frac{1}{L}\sum_{k\neq 0}e^{-ikR}\Big{]},$ (79a)
$g_{1}(R,t)=n(t).$ (79b)
In the thermodynamic limit, i.e. $L\rightarrow+\infty$, and since $R\in\mathbb{N}$, we have $(1/L)\sum_{k\neq 0}e^{-ikR}=\delta_{R,0}$ and thus $g_{2}$ takes the simple form $g_{2}(R,t)=n(t)[n(t)+\delta_{R,0}]$. By summing over the momentum $k$ the EoM associated to $G_{k}(t)=\langle\hat{n}_{k}\rangle_{t}$ the time-dependent occupation number in reciprocal space at Eq. (3), while considering the 1D case as well as on-site two-body losses, the EoM associated to the time-dependent density $n(t)$ the total occupation number can be deduced despresnew and it yields:
$\frac{\mathrm{d}}{\mathrm{d}t}n(t)=-\frac{2\gamma}{L}\sum_{R}\langle\hat{b}^{{\dagger}}_{R}\hat{b}^{{\dagger}}_{R}\hat{b}_{R}\hat{b}_{R}\rangle_{t}.$ (80)
Starting from the latter expression and using the canonical commutation rules as well as the translational invariance of the BH model together with Eq. (79), we obtain:
$\frac{\mathrm{d}}{\mathrm{d}t}n(t)=-2\gamma\left[g_{2}(0,t)-g_{1}(0,t)\right]=-2\gamma n(t)^{2},$ (81)
which can be solved analytically and we finally recover the result from Ref. despresnew regarding the condensate density $n(t)$ for the non-interacting case, i.e. $U=0$, namely:
$n(t)=\frac{\bar{n}}{1+2\gamma\bar{n}t},$ (82)
where $\bar{n}=\bar{n}(0)$ refers to the initial filling of the lattice chain. We stress that the previous theoretical form for the condensate density $n(t)$ is valid provided that the thermodynamic limit as well as small observation times are considered. Note that a more generic analytical expression for the latter quantity is provided at Ref. despresnew .